An investment with a 3year life and a cost of #120000 generates revenue of #25000 in year1 #45000 in year2 and #65000 in year3. If the discount rate is 8%

(i) what is the NPV of the investment.

(ii) state whether the investment should be accepted or not and why.

(Using the 12% as your second discount rate to solve the remaining part of the question under IRR.


1
Expert's answer
2021-08-02T11:04:06-0400

i)NPV = Present value of cash inflow - present value of cash outflow.

Or

"NPV = \frac{cash\space flow }{(1+i)^t} – initial\space investment"

Where,

i= discount rate

t= number of time period


"NPV = \frac{sum \space of \space cash\space flow }{(1+i)^t} – initial\space investment" "NPV = 23145.15+38580.25+51599.10-120000"

= -6672.51

NPV of the project is -$6672.51


ii.

Net present value of the project should be at least positive to be accepted for the project. As in this case, the NPV is -$6672.50, the project should be rejected at this discount rate and cash flow.